OpenStudy (anonymous):

What is the area of a rectangle with a length of 15 feet and a width of 9 feet?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \text{Area of a rectangle = Length } \times \text{Width} }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \text{Area of this rectangle =} 15 \times 9 }\)

OpenStudy (anonymous):

15*9 = 135 sq.ft

OpenStudy (anonymous):

area=l*b=15*9=135

OpenStudy (anonymous):

Thanks, i thought that was it.